OpenStudy (anonymous):

Anyone here go to James Madison, can maybe be a life saver and help me out with geometry 2? :)

OpenStudy (aravindg):

do you have a particular question? Perhaps we can help....

OpenStudy (anonymous):

I am waiting to see the problem as well..

OpenStudy (anonymous):

Lol. Find the area of a sector with a central angle of 170 and a diameter of 9.1 cm. Round to the nearest tenth. Options are A. 122.9 \[cm ^{2}\] B. 30.7 \[cm ^{2}\] C. 8.6 \[cm ^{2}\] D. 3.4 \[cm ^{2}\]

OpenStudy (aravindg):

thats better :)

OpenStudy (aravindg):

do you know formula for area of sector ?

OpenStudy (aravindg):

@LyssaLauren ?

OpenStudy (anonymous):

Before you do ANYTHING, convert that degree measure to radians.

OpenStudy (anonymous):

well if that is case than would be this be a triangle that we are dealing with? I know this isn't trig

OpenStudy (anonymous):

@LyssaLauren

OpenStudy (anonymous):

Yes but we have a formula for a sector.

OpenStudy (anonymous):

dido525 do you the shape is triangle?

OpenStudy (anonymous):

We must have that angle in degrees for that formula to work,

OpenStudy (aravindg):

@Dido525 not necessary here AREA OF SECTOR=\\large \dfrac{\theta}{360} \pi r^2\)

OpenStudy (anonymous):

No no, I know it's going to look like a slice of pizza lol.

OpenStudy (aravindg):

AREA OF SECTOR=\(\large \dfrac{\theta}{360} \pi r^2\)

OpenStudy (anonymous):

Te area of a sector of a circle is the product of the ratio measure of the arc 360 and the area of the circle. Right?

OpenStudy (aravindg):

theta is in degrees

OpenStudy (anonymous):

Ohh really? I didn;t know about that formula. Normally I just use A=A*r*Theta (in radians)

OpenStudy (anonymous):

But that formula makes perfect sence too.

OpenStudy (anonymous):

A=r*Theta(in radians) sorry. Typo.

OpenStudy (anonymous):

no prob man

OpenStudy (aravindg):

The A you are speaking of is arc length @Dido525

OpenStudy (aravindg):

we are finding area of sector here which is found by the formula I gave above

OpenStudy (anonymous):

Yeah. Stupid moment.

OpenStudy (aravindg):

@LyssaLauren did you get the answer?

OpenStudy (anonymous):

Sorry about that.

OpenStudy (aravindg):

np ...evryone makes mistakes ..next time be a bit more alert ! :)

OpenStudy (anonymous):

It would be B, right?

OpenStudy (anonymous):

Dido525 where I come from that is called a brain fart lol

OpenStudy (aravindg):

wait a min pls

OpenStudy (aravindg):

yes you are right!!

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

Thank you for your help! :]

OpenStudy (aravindg):

yw :)

OpenStudy (anonymous):

Nice job all..